OpenStudy (anonymous):

differentiate f(x)=x/(1+x^2)

OpenStudy (anonymous):

\[ f'(x)=\frac {1+x^2 -2 x x}{(1+x^2)^2}=\frac {1-x^2}{(1+x^2)^2} \]

OpenStudy (anonymous):

Apply quotient's rule.

OpenStudy (anonymous):

could you show me how?

OpenStudy (anonymous):

\[ \frac{u'v - v'u }{v^2} \]

OpenStudy (anonymous):

\[ u= x\\ v=1+x^2 \]

OpenStudy (anonymous):

\[ u'=1\\ v'= 2 x \]

OpenStudy (anonymous):

Can you take it from here?

OpenStudy (anonymous):

ahh yes i have figured it out :) thank you :) do you know if there a trick for knowing what rule to use?

OpenStudy (anonymous):

if it's a quotient, use quotient rule. If it's product, us product's rule